OpenStudy (hanner_b_nanner):

Fan&&Medal..Someone check my answer please! Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)). >>-40x + 24<< -40x + 31 -40x + 64 -40x + 70

OpenStudy (bradely):

f(g(x)) =f(10x - 6.) =-4(10x - 6)+7 =-40x+24+7 =-40x+31 Option 2 is correct Source: http://www.mathskey.com/question2answer/

OpenStudy (hanner_b_nanner):

ohhhhh, ok...thank you!